OpenStudy (ttp):

Jacob and Roman are standing next to each other. Jacob, who is one hundred eighty centimeters tall, has a shadow that is thirty centimeters long. If Roman is one hundred sixty-eight centimeters tall, approximately how long is his shadow? six centimeters eighteen centimeters twenty-eight centimeters forty-two centimeters

OpenStudy (anonymous):

 The two triangles are similar: \[\frac{ 180 }{ 168 } = \frac{ 30 }{ x }\] Therefore \[x = \frac{ 168*30 }{ 180 } = 28\] So the third answer.